# Relativity with or without light and Maxwell

### v2

@@ 1 Introduction
Recent papers published in pedagogical journals [1-9] testify that the problem of how to teach special relativity continues to be a fascinating topic, discussed at various levels of sophistication. This is no wonder because the concepts of where, when and why are essential for our existence as conscious beings, and with the theory of relativity our instinctive world view is at stake. It is a common experience of high school physics teachers that even the students with aversion to physics lose their apathy and listen with ears wide open when confronted with the theory of relativity .
The historical path to special relativity starts from the second postulate introduced by Einstein in 1905 [10]. Immediately after the publication of reference [10], he noted that the second postulate is contained in Maxwell’s equations [11], suggesting thus the approach recently elaborated by Aguirregabiria et al ([1], [12]).
In the present paper, which was inspired by references [1-12], we venture to illuminate some obscure corners and to simplify some involved points in the relativistic arguments. We believe that the present attempt could save some time and effort to the teachers and students of special relativity at undergraduate level, and provides some fresh insights. In Section 2 we discuss the complex relationship between Maxwell’s equations and the second postulate. Section 3 is an attempt to deduce the main results of the ‘relativity without light’ approach in a simple way. The concluding Section 4 contains a praise of the second postulate and a historical remark on the peculiar status of the principle of relativity among the Maxwellians in the late nineteenth century.
@@ 2 Maxwell’s equations, ‘inertia-time’ and ‘light-time’
Starting from the postulated validity of Maxwell’s equations in any given inertial frame involves that the meaning of ‘time’ has already been settled in that frame, according to the standard definition based on Galileo’s principle of inertia (‘inertia-time’). As is well known, inertia-time and the validity of Maxwell’s equations in any one inertial frame $S$ imply that one can also introduce ‘light-time’ in that frame, based on rectilinear propagation of electromagnetic waves [^1] Following Maxwell (1864), “light itself […] is an electromagnetic disturbance in the form of waves propagated through the electromagnetic field according to electromagnetic laws.” [13] in vacuum at the same speed $c\equiv(\epsilon_{0}\mu_{0})^{-1/2}$ in all directions; light-time is identical with inertia-time in $S$. Applying now the principle of relativity (assuming also that $\epsilon_{0}$ and $\mu_{0}$ are frame-independent), it follows that one should look for coordinate transformations between two inertial frames $S$ and $S^{\prime}$, in uniform motion with respect to one another, that are consistent with light-time. (Inertia-time and light-time are identical in any frame, and inertia-time puts no constraints on the transformations sought, except that the validity of Galileo’s principle of inertia in $S$ should imply its validity in $S^{\prime}$.) Clearly, the transformations cannot be of the Galilean type, involving $t^{\prime}=t$, since such transformations destroy invariance of $c$ [^2] The fact that ‘now’ is not the same for all inertial observers is a miracle of different kind than the one that the motion of the moon about the earth and the falling of an apple have a common root. (Newton proved the latter in the third book of Principia.) According to Newton, “relative, apparent and common time is any sensible and external measure (whether accurate or unequable) of duration by means of motion;” “duration or the perseverance of the existence of things” is Newton’s synonym for his “absolute, true and mathematical time.” Einstein’s postulated uniform propagation of light in vacuum, an ideal time–keeper (Silberstein’s term) identical for all inertial observers and inaccessible to experimental verification, appears to be a perfect analogon of Newton’s absolute time which “of itself and of its own nature, without relation to anything external, flows equably,” and the parts of which “make no impressions on the senses” (“non incurrunt in sensus;” all quotations of Newton are from the first Scholium of Principia [14]). Newton, and all physicists before Einstein (including Voigt, Larmor, Lorentz and Poincaré [15-18]), took it for granted that there was only one ‘time,’ absolute Newtonian time, for all observers in motion with respect to one another. Einstein was bold enough to venture that each inertial observer has her/his own absolute Einsteinian time. Assumed uniformity of space and time implies that the transformations must be linear; employing also invariance of the speed of light $c$ and assuming isotropy of space, one obtains the Lorentz transformations. [^3] Einstein’s original derivation of the Lorentz transformations [10], while cumbersome, is perfectly correct, without involving Galilean transformations, as clarified by Martinez [19]. Einstein for the first time gave the explanation of how uniformity of space and time implies the linearity of the transformations in [20], Section 7.4; in the same paper he gave for the first time a definition of a clock ([20], p 21).
@@
However, as noted in [22], Einstein in 1905 would have had to derive the Lorentz transformations following the template of thermodynamics, making them independent of the Maxwell theory, even if he were quite ignorant of light quanta or Planck’s 1900 derivation of the black–body radiation formula. A definition of time that is as simple as possible must conceptually precede any discussion about “the laws according to which the states of physical systems change,” including the domain of the validity of Maxwell’s equations. Attempting to keep the Lorentz transformations divorced from the Maxwell electromagnetic theory, Einstein took a necessary condition of the Maxwell theory to be his principle of the constancy of the speed of light (the notorious “second postulate”). This principle, which appears to be ‘simplicity itself,’ combined with the (meta-)principle of relativity, is the crux of special relativity. Einstein’s ‘definite velocity $V$’ of propagation of light in vacuum (he used the symbol ‘$V$’ instead of ‘$c$’ in [10]), appearing in the introduction and in paragraph 2 of [10], does not have the familiar meaning of a derived quantity, $V=$ (light path)/time interval, where ‘time interval’ is a segment of inertia–time. Instead, Einstein’s $V$ is a primitive quantity (of course, all other speeds are derived quantities), and ‘time interval’ is by definition equal to (one–way light path)/$V$. Thus the speed of light in vacuum is by definition equal to the universal constant $V$, identical for all inertial observers, in perfect disagreement with our instinctive Galilean mentality. [^4] As is recalled in [23], Einstein’s second postulate has a rather intricate content. “It postulates that, relative to any given inertial frame, the one-clock two-way speed of light in vacuum $V$ (a measurable physical quantity and, as measurements reveal, a universal constant) is constant and independent of the velocity of light source and equals the one-way two-clock speed of light in vacuum (an immeasurable quantity). […] It should be pointed out that throughout the Relativity Paper Einstein used the same symbol (‘$V$’) for the speed of light in vacuum and the phase velocity of electromagnetic waves in vacuum according to the ‘Maxwell-Hertz equations’, linking thus special relativity with Maxwell’s theory, and at the same time linking the new time keeper (propagation of light) with the earlier ones (cf [24]).” Two circumstances were decisive for Einstein’s argument: first, Maxwell’s equations can be simply made Lorentz–covariant and thus consistent with the principle of relativity by appropriately defining the transformations for the electric and magnetic fields and charge and current densities (the fact already known (partly) to Lorentz [25] and (fully) to Poincaré [26], but also (partly) to Larmor [27], cf [16]), and, second, empirical evidence for Maxwell’s theory in pseudo–inertial frame tied to the earth.
@@ 3 Relativity without the light postulate
1. There is a frame of reference in which spatial coordinates $x$, $y$ and $z$ are Cartesian coordinates (Euclidean geometry applies), and time coordinate $t$ is defined through ‘light–time’ (Einstein’s second postulate applies so the one–way two–clock speed of light is by definition the known constant, $c\equiv(\epsilon_{0}\mu_{0})^{-1/2}$); we will call a frame of reference with such properties of space and time the ‘primitive frame.’
@@
where $F_{v}$ and $G_{v}$ are parameters dependent solely on $v$, as yet unknown. Setting for simplicity $y^{\prime}=y$ and $z^{\prime}=z$, our considerations will be restricted to $x$ and $t$ only. Employing the principle of relativity, assuming velocity reciprocity [^5] Velocity reciprocity means that the velocity of the $S^{\prime}$ frame with respect to the $S$ frame is the opposite of the velocity of $S$ with respect to $S^{\prime}$. Many an author, including Ignatowski [28], takes velocity reciprocity as an immediate and self-evident consequence only of the principle of relativity. However, various additional assumptions are also required, such as uniformity of space and of time, spatial and temporal isotropy, and causality, as pointed out by Berzi and Gorini [29] and by Lévy-Leblond [30]; a thorough discussion of the issue was recently published by Patrick Moylan [5]. By the way, Einstein’s first derivation of the Lorentz transformations (given in paragraph 3 of [10]) based on the postulated, finite and known, universal speed $V$ (Einstein’s symbol for ‘$c$’), does not assume velocity reciprocity, but deduces it. It is perhaps amusing to note that therein he denotes by $\varphi(v)$ both his $a$ and $a\beta$, where $\beta\equiv\frac{1}{\sqrt{1-(v/V)^{2}}}$. Moreover, he employs three reference frames (which he calls ‘coordinate systems’), ‘resting frame’ $K$ with coordinates $x,y,z$ and $t$, ‘moving frame’ $k$ with coordinates $\xi,\eta,\zeta$ and $\tau$, and a third frame $K^{\prime}$ with coordinates $x^{\prime},y^{\prime},z^{\prime}$ and $t^{\prime}$, moving relative to $k$. In the same paragraph 3 of [10] he uses the symbol $x^{\prime}$ also for $x-vt$, so $x^{\prime},y,z$ and $t$ refer to coordinate systems introduced in the reference frame $K$. All that of course does not add to lucidity; reading Einstein, like reading Maxwell, is always an adventure par excellence [31,32]., one has
@@
which excludes the possibility of negative $\Omega$ (otherwise, one would have $01+\Omega uv=$ 0 for $uv=1/|\Omega|$). Thus the universal constant $\Omega$ must be nonnegative,
@@
which imply that there is one time for all inertial observers (identical with Newton’s absolute time), that a standard of length always has the same length independent of its velocity, and that $v$ can be arbitrarily large. While in perfect agreement with our instinctive Galilean mentality, the Galilean transformations cannot be made compatible with a consistent interpretation of experience. [^6] Newton’s absolute space (that “by its own nature without relation to any thing external remains always uniform and unchangeable”) and absolute time (that “by its own nature without relation to any thing external flows equably”) appear to be the natural habitat for the Galilean transformations. As the parts of absolute space “do by no means come under the observation of our senses” (“non incurrunt in sensus”), and the same applies to absolute time, it appears that those absolute concepts remain transcendental and thus perhaps fruitless (compare Maxwell’s poetic discussion in his Matter and Motion [36]; by the way, what Maxwell calls “the doctrine of relativity of all physical phenomena” is the root of Poincaré–Einstein’s principle of relativity). But, keeping in mind that “everything which is not forbidden is allowed,” one should not dispense with a possible interpretation of the Galilean transformations.
@@
Thus $\overline{c}$ is the universal limit speed for massive particles (a supremum of possible particle speeds), as yet unknown, for the case $\Omega>0$. Putting also
@@
Since we concluded above that $w$ is the speed of $S^{\prime\prime}$ relative to $S$, in eq. (25) one recognizes the familiar relativistic composition of velocities, and in eq. (26) the familiar relativistic transformation for gamma factors (cf, e. g., [37, 38]).
@@ 4 Concluding comments
Turning to phenomena, however, is tricky considering that inertial observers inhabit “ideal infinitely extended gravity-free inertial frames,” whereas in the real world their habitat is restricted to “the freely falling nonrotating local frames” [37]. [^7] This fact was basically well known to Newton, see corollaries V and VI of the Laws of Motion [14] and also [39]. We quoted above Einstein’s clear-cut statement that “the Lorentz transformation, the real basis of the special relativity theory, in itself has nothing to do with the Maxwell theory […],” which is widely endorsed in the literature. (“Special relativity would exist even if light and electromagnetism were somehow eliminated from nature.”[37]) While various scenarios that have been proposed to ascertain the value of $\overline{c}$, without invoking properties of light or Maxwell’s electromagnetic theory, appear to be free from logical errors (cf, e. g., [6, 33, 40]), none of them can be actually implemented.
@@
On the other hand, Einstein’s second postulate combined with the principle of relativity states, basically: first, there is a universal finite speed, c, the same for all primitive observers, and light propagates in vacuum at the universal speed; second, the universal speed c is a primitive quantity, and ‘time’ (‘light–time’, ‘duration’) in any primitive frame is a derived quantity through time interval =$\stackrel{{\scriptstyle\mbox{\small d}}}{{=}}$ light path/c; third, the magnitude of $c$ is equal to the speed of propagation of electromagnetic waves in vacuum, $(\epsilon_{0}\mu_{0})^{-1/2}$ as given by Maxwell’s electromagnetic theory, and also it is equal to the one-clock two-way speed of light as ascertained by terrestrial measurements and consistent with astronomical observations.
@@
Thus, whereas in ‘relativity without light and electromagnetism’ $\overline{c}$ is the universal limit speed determinable only in principle through Gedankenexperiments, in the second postulate approach $c$ is the speed which pertains to a real phenomenon (the essential one for our experience of the world), “concerning which we know something certain […] in a higher degree than for any other process which could be considered, thanks to the investigations of Maxwell and H. A. Lorentz” [41]. [^8] It should be stressed that relativistic length contraction and clock retardation cannot be verified directly. As Jefimenko [42] notes, Einstein’s method for measuring the length of a moving rod proposed in [10], “was, of course, merely a ‘Gedankenexperiment,’ that is, an imaginary procedure, a verbalization of [equation $x^{\prime}=\gamma(x-vt)$], that cannot be actually implemented.” The second postulate combined with the principle of relativity, while counterintuitive, gives a simple path to the concepts of time and primitive frame, and to the Lorentz transformations. The second postulate is more fundamental than the principle of relativity because a definition of time that is as simple as possible must conceptually precede any discussion about “the laws under which the states of physical systems undergo changes.”
@@
To summarize, Einstein’s definition of the “time” [10] is basically based on a tacit assumption that there is a universal finite speed of propagation of interaction; the limit speed is a primitive quantity equal to the speed of light from Maxwell’s electromagnetic theory, $c$, and “time interval” is a derived quantity. On the other hand, in Ignatowski’s approach, a universal finite speed of propagation of interaction, $\overline{c}$, arises as a consequence of his starting assumptions; the value of the limit speed $\overline{c}$ remains unknown, perhaps transcendent, independent of any physical theory. In this sense Ignatowski’s relativity is more general than Eistein’s relativity [43], true, the meaning of Ignatowski’s “$t$” is somewhat blurred.

### v3

@@ Abstract
keywords:
special relativity, the Maxwell electromagnetic theory, time, light, the second postulate, Ignatowski’s relativity, the principle of relativity
@@ 1 Introduction
Recent papers published in pedagogical journals [1-9] testify that the problem of how to teach special relativity continues to be a fascinating topic, discussed at various levels of sophistication. This is no wonder because the concepts of where, when and why are essential for our existence as conscious beings, and with the theory of relativity our instinctive world view is at stake. It is a common experience of high school physics teachers that even the students with an aversion to physics lose their apathy and listen with ears wide open when confronted with the theory of relativity .
The historical path to special relativity stricto sensu began with the second postulate introduced by Einstein in 1905 [10]. Immediately after the publication of reference [10], he noted that the second postulate is contained in Maxwell’s equations [11], suggesting thus the approach recently elaborated by Aguirregabiria et al ([1], [12]).
In the present paper, which was inspired by references [1-12], we venture to illuminate some dark corners and to simplify some involved points in relativistic arguments. We believe that the present attempt could save some time and effort to the teachers and students of special relativity at undergraduate level, and provides some fresh insights. In Section 2 we discuss the complex relationship between Maxwell’s equations and the second postulate. Section 3 is an attempt to deduce the main results of the ‘relativity without light’ approach in a simple way. The concluding Section 4 contains a praise of the second postulate and a historical remark on the peculiar status of the principle of relativity among the Maxwellians in the late nineteenth century.
@@ 2 Maxwell’s equations, ‘inertia-time’ and ‘light-time’
Starting from the postulated validity of Maxwell’s equations in an inertial frame $S$ involves that the meaning of ‘time’ has already been settled in the frame $S$, according to the standard definition based on Galileo’s principle of inertia (‘inertia-time’). As is well known, inertia-time and the validity of Maxwell’s equations in the frame $S$ imply that one can also introduce ‘light-time’ in that frame, based on rectilinear propagation of electromagnetic waves [^1] Following Maxwell (1864), “light itself […] is an electromagnetic disturbance in the form of waves propagated through the electromagnetic field according to electromagnetic laws.” [13] in vacuum at the same speed $c\equiv(\epsilon_{0}\mu_{0})^{-1/2}$ in all directions; light-time is identical with inertia-time in $S$. Applying now the principle of relativity (assuming also that $\epsilon_{0}$ and $\mu_{0}$ are frame-independent), it follows that one should look for coordinate transformations between two inertial frames $S$ and $S^{\prime}$, in uniform motion with respect to one another, that are consistent with light-time. (While inertia-time and light-time are identical in any frame, inertia-time puts no constraints on the transformations sought except that the validity of Galileo’s principle of inertia in $S$ should imply its validity in $S^{\prime}$.) Clearly, the transformations cannot be of the Galilean type, involving $t^{\prime}=t$, since such transformations destroy invariance of cc. [^2] The fact that ‘now’ is not the same for all inertial observers is a miracle of different kind than the one that the motion of the moon about the earth and the falling of an apple have a common root. (Newton proved the latter in the third book of Principia.) According to Newton, “relative, apparent and common time is any sensible and external measure (whether accurate or unequable) of duration by means of motion;” “duration or the perseverance of the existence of things” is Newton’s synonym for his “absolute, true and mathematical time.” Einstein’s postulated uniform propagation of light in vacuum, an ideal time–keeper (Silberstein’s term) identical for all inertial observers and inaccessible to experimental verification, appears to be a perfect analogon of Newton’s absolute time which “of itself and of its own nature, without relation to anything external, flows equably,” and the parts of which “make no impressions on the senses” (“non incurrunt in sensus;” all quotations of Newton are from the first Scholium of Principia [14]). Newton, and all physicists before Einstein (including Voigt, Larmor, Lorentz and Poincaré [15]-[18]), took it for granted that there was only one ‘time,’ absolute Newtonian time, for all observers in motion with respect to one another. Einstein was bold enough to venture that each inertial observer has her/his own absolute Einsteinian time. Assumed uniformity of space and time implies that the transformations must be linear; employing also invariance of the speed of light $c$ and assuming isotropy of space, one obtains the Lorentz transformations. [^3] Einstein’s original derivation of the Lorentz transformations [10], while cumbersome, is perfectly correct, without involving Galilean transformations, as clarified by Martínez [19]. Einstein for the first time gave the explanation of how uniformity of space and time implies the linearity of the transformations in [20], Section 7.4; in the same paper he gave for the first time a definition of a clock ([20], p 21), that involves “the principle of sufficient reason.”
@@
However, as noted in [22], Einstein in 1905 would have had to derive the Lorentz transformations following the template of thermodynamics, making them independent of the Maxwell theory, even if he had been completely ignorant of light quanta or Planck’s 1900 derivation of the black–body radiation formula. Namely, a definition of time that is as simple as possible must conceptually precede any discussion about “the laws according to which the states of physical systems change” (“die Gesetze nach denen sich die Zustände der physikalischen Systeme ändern” [10]), including the domain of the validity of Maxwell’s equations. Attempting to keep the Lorentz transformations divorced from the Maxwell electromagnetic theory, Einstein took a necessary condition of the Maxwell theory, replacing “electromagnetic waves” with “light,” to be his principle of the constancy of the speed of light (the notorious “second postulate”). This principle, which appears to be ‘simplicity itself,’ combined with the (meta-)principle of relativity, is the crux of special relativity. Einstein’s ‘definite velocity $V$’ of propagation of light in vacuum (he used the symbol ‘$V$’ instead of ‘$c$’ in [10]), appearing in the introduction and in paragraph 2 of [10], does not have the familiar meaning of a derived quantity, $V=$ (light path)/time interval, where ‘time interval’ is a segment of inertia–time. Instead, Einstein’s $V$ is a primitive quantity (of course, all other speeds are derived quantities), and ‘time interval’ is by definition equal to (one–way light path)/$V$. Thus the speed of light in vacuum is by definition equal to the universal constant $V$, identical for all inertial observers, in perfect disagreement with our instinctive Galilean mentality. [^4] As is recalled in [23], Einstein’s second postulate has a rather intricate content. It postulates that, relative to an inertial frame, “the one-clock two-way speed of light in vacuum $V$ (a measurable physical quantity and, as measurements reveal, a universal constant) is constant and independent of the velocity of light source and equals the one-way two-clock speed of light in vacuum (an immeasurable quantity). […] It should be pointed out that throughout the Relativity Paper Einstein used the same symbol (‘$V$’) for the speed of light in vacuum and the phase velocity of electromagnetic waves in vacuum according to the ‘Maxwell-Hertz equations’, linking thus special relativity with Maxwell’s theory, and at the same time linking the new time keeper (propagation of light) with the earlier ones (cf [24]).” Two circumstances were decisive for Einstein’s argument: first, Maxwell’s equations can be simply made Lorentz–covariant and thus consistent with the principle of relativity by appropriately defining the transformations for the electric and magnetic fields and charge and current densities (the fact already known (partly) to Lorentz [25] and (fully) to Poincaré [26], but also (partly) to Larmor [27], cf [16]), and, second, empirical evidence for Maxwell’s theory in pseudo–inertial frame tied to the earth.
@@ 3 Relativity without the light postulate
1. There exists a frame of reference in which spatial coordinates $x$, $y$ and $z$ are Cartesian coordinates (Euclidean geometry applies), and time coordinate $t$ is defined through ‘light–time’ (light signals in vacuum propagate rectilinearly and uniformly and the one–way two–clock speed of light is by definition the known constant, $c\equiv(\epsilon_{0}\mu_{0})^{-1/2}$); we will call a frame of reference with such properties of space and time the ‘primitive frame.’
@@
where $F_{v}$ and $G_{v}$ are parameters dependent solely on $v$, as yet unknown. Setting for simplicity $y^{\prime}=y$ and $z^{\prime}=z$ (for a convincing deduction see [10]), our considerations will be restricted to $x$ and $t$ only. Employing the principle of relativity, assuming velocity reciprocity [^5] Velocity reciprocity means that the velocity of the $S^{\prime}$ frame with respect to the $S$ frame is the opposite of the velocity of $S$ with respect to $S^{\prime}$. Many an author, including Ignatowski [28], takes velocity reciprocity as an immediate and self-evident consequence only of the principle of relativity. However, various additional assumptions are also required, such as uniformity of space and of time, spatial and temporal isotropy, and causality, as pointed out by Berzi and Gorini [29] and by Lévy-Leblond [30]; a thorough discussion of the issue was recently published by Patrick Moylan [5]. By the way, Einstein’s first derivation of the Lorentz transformations (given in paragraph 3 of [10]) based on the postulated, finite and known, universal speed $V$ (Einstein’s symbol for ‘$c$’), does not assume velocity reciprocity, but deduces it. It is perhaps amusing to note that therein he denotes by $\varphi(v)$ both his $a$ and $a\beta$, where $\beta\equiv\frac{1}{\sqrt{1-(v/V)^{2}}}$. Moreover, he employs three reference frames (which he calls ‘coordinate systems’), ‘resting frame’ $K$ with coordinates $x,y,z$ and $t$, ‘moving frame’ $k$ with coordinates $\xi,\eta,\zeta$ and $\tau$, and a third frame $K^{\prime}$ with coordinates $x^{\prime},y^{\prime},z^{\prime}$ and $t^{\prime}$, moving relative to $k$. In the same paragraph 3 of [10] he uses the symbol $x^{\prime}$ also for xvtx-vt, so $x^{\prime},y,z$ and $t$ refer to coordinate systems introduced in the reference frame $K$. All that of course does not add to lucidity; reading Einstein, like reading Maxwell, is always an adventure par excellence [31,32]., one has
@@
which excludes the possibility of negative $\Omega$ (otherwise, one would have $1+\Omega uv=0$ for $uv=1/|\Omega|$). Thus the universal constant $\Omega$ must be nonnegative,
@@
which imply that there is one time for all inertial observers (identical with Newton’s absolute time), that a standard of length always has the same length independent of its velocity, and that $v$ can be arbitrarily large. While in perfect agreement with our instinctive Galilean mentality, the Galilean transformations cannot be made compatible with a consistent interpretation of experience. [^6] Newton’s absolute space (that “by its own nature without relation to any thing external remains always uniform and unchangeable”) and absolute time (that “by its own nature without relation to any thing external flows equably”) appear to be the natural habitat for the Galilean transformations. As the parts of absolute space “do by no means come under the observation of our senses” (“non incurrunt in sensus”), and the same applies to absolute time, it appears that those absolute concepts remain transcendental, in Kantian sense (compare Maxwell’s poetic discussion in his Matter and Motion [36]; by the way, what Maxwell calls “the doctrine of relativity of all physical phenomena” is the root of Poincaré–Einstein’s principle of relativity). But, keeping in mind that “everything which is not forbidden is allowed,” one should not dispense with a possible interpretation of the Galilean transformations. It should be stressed that in special relativity all inertial frames share one and the same space, cf. Section 3 of [10], which is analogous to Newton’s absolute space, without of course involving the idea of absolute rest.
@@
Thus $\overline{c}$ is the universal limit speed for massive particles, as yet unknown. Putting also
@@
Since we concluded above that $w$ is the speed of $S^{\prime\prime}$ relative to $S$, in eq. (25) one recognizes the familiar relativistic composition of velocities, and in eq. (26) the familiar relativistic transformation for gamma factors (cf, e. g., [37, 38]). It can be shown, employing eqs. (24) and invariance of causality, that the speed of propagation of interaction must be less than or equal to $\overline{c}$, but the theory does not specify whether $\overline{c}$ is a physically attainable speed or not [37].
@@ 4 Concluding comments
Turning to phenomena, however, is tricky considering that inertial observers inhabit “ideal infinitely extended gravity-free inertial frames,” whereas in the real world their habitat is restricted to “the freely falling nonrotating local frames” [37]. [^7] This fact was basically well known to Newton, see corollaries V and VI of the Laws of Motion [14], and also [39]. We quoted above Einstein’s clear-cut statement that “the Lorentz transformation, the real basis of the special relativity theory, in itself has nothing to do with the Maxwell theory […],” which is widely endorsed in the literature. (“Special relativity would exist even if light and electromagnetism were somehow eliminated from nature.”[37]) While various scenarios that have been proposed to ascertain the value of $\overline{c}$, without invoking properties of light or Maxwell’s electromagnetic theory, appear to be free from logical errors (cf, e. g., [6, 33, 40]), none of them can be actually implemented.
@@
On the other hand, Einstein’s second postulate combined with the principle of relativity states, basically: first, there is a universal finite speed, c, the same for all primitive observers, and light propagates in vacuum at the universal speed; second, the universal speed c is a primitive quantity, and ‘time’ (‘light–time’, ‘duration’) in any primitive frame is a derived quantity through time interval $\stackrel{{\scriptstyle\mbox{\small d}}}{{=}}$ light path/c; third, the magnitude of $c$ is equal to the speed of propagation of electromagnetic waves in vacuum, $(\epsilon_{0}\mu_{0})^{-1/2}$ as given by Maxwell’s electromagnetic theory, and also it is equal to the one-clock two-way speed of light as ascertained by terrestrial measurements and consistent with astronomical observations.
@@
Thus, whereas in ‘relativity without light and electromagnetism’ $\overline{c}$ is the universal limit speed determinable only in principle through Gedankenexperiments, in the second postulate approach $c$ is the speed which pertains to a real phenomenon (the essential one for our experience of the world), “concerning which we know something certain […] in a higher degree than for any other process which could be considered, thanks to the investigations of Maxwell and H. A. Lorentz” [41]. The second postulate combined with the principle of relativity, while counterintuitive, gives a simple path to the concepts of time and primitive frame, and to the Lorentz transformations. [^8] It should be stressed that relativistic length contraction and clock retardation cannot be verified directly. As Jefimenko [42] notes, Einstein’s method for measuring the length of a moving rod proposed in [10], “was, of course, merely a ‘Gedankenexperiment,’ that is, an imaginary procedure, a verbalization of [equation $x^{\prime}=\gamma(x-vt)$], that cannot be actually implemented.” Also, some familiar relativistic generalizations concerning clock retardation have to be amended [43]. The second postulate is more fundamental than the principle of relativity because a definition of time that is as simple as possible must conceptually precede any discussion about “the laws according to which the states of physical systems change.”
@@
To summarize, Einstein’s definition of the “time” [10] is basically based on a tacit assumption that there is a universal finite speed of propagation of interaction; the limit speed is a primitive quantity equal to the speed of light from Maxwell’s electromagnetic theory, $c$, and “time interval” is a derived quantity. On the other hand, in Ignatowski’s approach, the universal limit speed of propagation of interaction (whether attainable or not), $\overline{c}$, arises as a consequence of his starting assumptions; the value of the limit speed $\overline{c}$ remains unknown, perhaps transcendent, independent of any physical theory. In this sense Ignatowski’s relativity is more general than Eistein’s relativity [44], true, the meaning of Ignatowski’s “tt” is somewhat blurred. Obviously, if $\overline{c}\neq c$, then Maxwell’s equations are not Lorentz-like covariant, and thus do not conform to Ignatowski’s relativity.
@@
Funding
This research received no external funding.
Data Availability Statement
No new data were generated or analyzed in support of this research.
Conflicts of interest
The author declares no conflicts of interest.